# RNA-FrameFlow: Flow Matching for de novo 3D RNA Backbone Design

### v1

@@ Abstract
We introduce RNA-FrameFlow, the first generative model for 3D RNA backbone design. We build upon $SE(3)$ flow matching for protein backbone generation and establish protocols for data preparation and evaluation to address unique challenges posed by RNA modeling. We formulate RNA structures as a set of rigid-body frames and associated loss functions which account for larger, more conformationally flexible RNA backbones (13 atoms per nucleotide) vs. proteins (4 atoms per residue). Toward tackling the lack of diversity in 3D RNA datasets, we explore training with structural clustering and cropping augmentations. Additionally, we define a suite of evaluation metrics to measure whether the generated RNA structures are globally self-consistent (via inverse folding followed by forward folding) and locally recover RNA-specific structural descriptors. The most performant version of RNA-FrameFlow generates locally realistic RNA backbones of 40-150 nucleotides, over 40% of which pass our validity criteria as measured by a self-consistency TM-score $\geq 0.45$, at which two RNAs have the same global fold.
@@ 1 Introduction
Designing RNA structures. Proteins, and the diverse structures they can adopt, drive essential biological functions in cells. Deep learning has led to breakthroughs in structural modeling and design of proteins (Jumper et al., 2021, Dauparas et al., 2022, Watson et al., 2023), driven by the abundance of 3D data from the Protein Data Bank (PDB). Concurrently, there has been a surge of interest in Ribonucleic Acids (RNA) and RNA-based therapeutics for gene editing, gene silencing, and vaccines (Doudna and Charpentier, 2014, Metkar et al., 2024). RNAs play a dual role as carriers of genetic information coding for proteins (mRNAs) as well as performing functions driven by their tertiary structural interactions (riboswitches and ribozymes). While there is growing interest in designing structured RNAs for a range of applications in biotechnology and medicine (Mulhbacher et al., 2010, Damase et al., 2021), the current toolkit for 3D RNA design uses classical algorithms and heuristics to assemble RNA motifs as building blocks (Han et al., 2017, Yesselman et al., 2019). However, hand-crafted heuristics are not always broadly applicable across multiple tasks and rigid motifs may not fully capture the conformational dynamics that govern RNA functionality (Ganser et al., 2019, Li et al., 2023a). This presents an opportunity for deep generative models to learn data-driven design principles from existing 3D RNA structures.
@@
Figure 1: The RNA-FrameFlow pipeline for 3D backbone generation. Our implementation establishes RNA-specific protocols for data preparation and evaluation for FrameFlow (Yim et al., 2023a). (1) Each nucleotide in the RNA backbone is converted into a frame to parameterize the placement of $C4^{\prime}$ by a translation, $C3^{\prime}-C4^{\prime}-O4^{\prime}$ by a rotation, and the rest of the atoms via 8 torsion angles $\Phi$. (2) We train generative models on all RNA structures of length 40-150 nucleotides from RNAsolo (Adamczyk et al., 2022). We also explore training with structural clustering and cropping augmentations to tackle the lack of diversity in 3D RNA datasets. (3) We introduce evaluation metrics to measure the recovery of local structural descriptors and global self-consistency of designed structures via inverse-folding with gRNAde (Joshi et al., 2023) followed by forward-folding with RhoFold (Shen et al., 2022).
@@
What makes deep learning for RNA hard? The primary challenge is the paucity of raw 3D RNA structural data, manifesting as an absence of ML-ready datasets for model development (Joshi et al., 2023). Protein structure is primarily driven by hydrogen bonding along the backbone, and current geometric deep learning models incorporate this inductive bias through backbone frames to represent residues (Jumper et al., 2021). RNA structure, however, is often more conformationally flexible and driven by base pairing interactions across strands as well as base stacking between rings of adjacent nucleotides (Vicens and Kieft, 2022), all of which can only be learnt implicitly at present111 See Eric Westhof’s talk contrasting RNA and protein structure. .
@@
Additionally, RNA nucleotides, the equivalent of amino acids in proteins, include significantly more atoms as part of the backbone (13 compared to 4) which necessitates a generalization of backbone frames where the placement of most atoms needs to be parameterized by torsion angles. These complexities have contributed to relatively poor performance of deep learning for RNA structure prediction compared to proteins (Kretsch et al., 2023, Abramson et al., 2024). Additionally, structure prediction models cannot directly be used for designing or generating novel RNA structures with desired constraints, which our work aims to do.
Our contributions. We develop RNA-FrameFlow, the first generative model for 3D RNA backbone design, illustrated in Figure 1. We adapt FrameFlow (Yim et al., 2023a), an $SE(3)$ equivariant flow matching model for proteins to RNA. We introduce RNA-specific modifications to the data preparation and loss formulation, including representing RNA nucleotides as rigid-body frames that parameterize all 13 atoms. We also introduce an evaluation pipeline to benchmark RNA backbone design models’ capabilities at recovering local and global structure. Our best model is trained on RNAs of lengths 40-150 from the PDB and can unconditionally sample locally plausible backbones with over 40% validity as measured by a self-consistency TM-score $\geq 0.45$.
Through this study, we aimed to evaluate the extent to which generative models for proteins can be adapted for RNA. This brought up critical challenges and limitations of deep learning for RNA modelling, such as a lack of explicit representations of the physical interactions that drive RNA structure as well as biases in 3D RNA datasets, which we have made preliminary efforts towards addressing. Together with our engineering contributions, we hope this work will stimulate future research in generative models for RNA design.
@@ 2 The RNA-FrameFlow Pipeline
Overview. We are concerned with building a generative model that unconditionally outputs all-atom RNA backbones, sampled from a distribution of realistic 3D RNA structures. Formally, given an RNA sequence length of $N_{\text{nt}}$ nucleotides, we aim to generate a real-valued tensor $\mathbf{X}$ of shape ${N_{\text{nt}}\times 13\times 3}$ representing 3D atomic coordinates for each of the 13 backbone atoms per nucleotide. In the following sections, we will describe how we adapt FrameFlow (Yim et al., 2023a), an $SE(3)$ equivariant flow matching model for protein backbones, to our setting.
@@ 2.1 Representing RNA Backbones as Frames
As shown in Figure 1, the RNA backbone consists of nucleotides with a phosphate group ($P,OP1,OP2,O5^{\prime}$), a ribose sugar ($C1^{\prime}-C5^{\prime},O2^{\prime},O3^{\prime},O4^{\prime}$), and a nitrogen atom $N$ at the stem of the base. We represent the group of atoms within each nucleotide as a rigid-body frame. Frames enable inferring the positions of all atoms within the nucleotide via a frame center and orientation (described subsequently). However, the 13 atoms per nucleotide in the RNA backbone is significantly greater than protein residues with 4 atoms ($C_{\alpha},N,C,O$). In proteins, it is standard to represent each residue by a frame centered at $C_{\alpha}$ with vectors along $C_{\alpha}-N$ and $C_{\alpha}-C$, and $O$ is placed assuming an idealised planar geometry (Jumper et al., 2021). No such canonical frame representation exists for RNAs.
RNA frames. We use the $C4^{\prime},C3^{\prime}$, and $O4^{\prime}$ atoms to create the frame for each nucleotide, as in Morehead et al. (2023). All other backbone atoms are inferred with 8 torsions $\Phi=\{\phi_{1}\rightarrow\phi_{8}\},\phi_{i}\in SO(2)$ that are predicted post-hoc after frame generation. The Gram-Schmidt process is used on $v_{1},v_{2}$ defined by the vectors along the $C4^{\prime}-O4^{\prime}$ and $C4^{\prime}-C3^{\prime}$ bonds; $C5^{\prime}$ is imputed based the positions of the other 3 atoms and tetrahedral geometry. Given the 8 torsion angles, we autoregressively place non-frame atoms in order of the torsions $\Phi$ in Figure 1, constructing the final all-atom RNA backbone structure.
Criteria on choosing frame atoms. We had two main considerations for selecting the subset of atoms to create RNA frames: (1) the atoms should have roughly the same spatial orientation w.r.t. each other; and (2) the atoms should be reasonably close to the centroid in the nucleotide to reduce error accumulation when placing the furthest non-frame atoms. We choose $C3^{\prime}$, $C4^{\prime}$, and $O4^{\prime}$ as these atoms spatially shift the least in naturally occurring RNA (Harvey and Prabhakaran, 1986). The non-frame backbone atoms – the remaining atoms in the ribose sugar ring and the phosphate group atoms – are parameterized by torsion angles to account for their relative conformational flexibility. This choice of frame enables models to learn ring puckering, the planar rotation of the ribose sugar ring about the $C4^{\prime}-C5^{\prime}$ bond which affects how the RNA interacts with partners to form complexes (Clay et al., 2017). We are actively evaluating alternate choices of RNA frame atoms.
@@ 2.2 $SE(3)$ Flow Matching on RNA Frames
Input. Given a set of 3D coordinates, a simultaneous rotation and translation $(r,x)$ forms an orientation-preserving rigid-body transformation of the coordinates. The set of all such transformations in 3D is the Special Euclidean group $SE(3)$, which composes the group of 3D rotations $SO(3)$ and 3D translations in $\mathbb{R}^{3}$.
We can represent an RNA frame $T=(r,x)$ as a translation $x\in\mathbb{R}^{3}$ from the global origin to place $C4^{\prime}$ and a rotation $r\in SO(3)$ to orient $C3^{\prime}-C4^{\prime}-O4^{\prime}$. Compared to working with raw 3D coordinates for each backbone atom, using the frame representation entails performing flow matching on the space of $SE(3)$. This is an inductive bias to reduce the degrees of freedom the generative model needs to learn. Instead of predicting 13 correlated 3D coordinates independently (39 quantities) for each nucleotide, we instead predict one 3D coordinate (of $C4^{\prime}$) and one 3$\times$3 rotation matrix (12 quantities). We follow Chen and Lipman (2024) and Yim et al. (2023a)’s framework for flow matching on $SE(3)$, which we summarise subsequently.
@@
Suppose $p_{0}(T_{0})$ and $p_{1}(T_{1})$ are the marginal distributions of randomly oriented and ground truth frames from our dataset of RNA structures, respectively. Suppose a non-unique time-dependent vector field $u_{t}$ leads to an ODE between the two distributions $p_{0}$ and $p_{1}$, i.e., assume there is a way to map from noisy samples to the corresponding true samples. This solution forms a ground truth probability path $p_{t}$ between the two distributions at time $t\in[0,1]$, which we can use to transform samples from noise to the true distribution. The continuity equation $\frac{\partial p}{\partial t}=-\nabla\cdot(p_{t}u_{t})$ relates the vector field $u_{t}$ to the evolution of the probability path $p_{t}$.
Given a noisy frame $T_{0}$ sampled from $p_{0}(T_{0})$ and the corresponding ground truth frame $T_{1}$ sampled from $p_{1}(T_{1})$, we construct a flow $T_{t}$ by following the probability path $p_{t}$ between $T_{0}$ and $T_{1}$ for any time step $t$ sampled from $\mathcal{U}(0,1)$. As shown by Chen and Lipman (2024) for the $SE(3)$ group (and other manifolds), the shortest path between the two states $T_{0}$ and $T_{1}$ can be used to define an interpolation:
@@
The architecture of the flow model $v_{\theta}$ is similar to the structure module from AlphaFold2 comprising Invariant Point Attention layers interleaved with standard Transformer encoder layers, following Yim et al. (2023a, b). We use an MLP head to predict torsion angles $\Phi$.
@@
Suppose $S=\{C4^{\prime},C3^{\prime},O4^{\prime}\}$ is the set of frame atoms [^2] In Section C.1, we show how including all backbone atoms better accounts for larger RNA nucleotides and improves validity of generated samples. and the sequence length is $N$. We summarise the auxiliary losses subsequently.
@@
Conditional generation. The unconditional sampling strategy described above aims to generate realistic RNA backbone structures sampled from the training distribution. However using generative models in real-world design tasks entails conditional generation based on specified design constraints or requirements (Ingraham et al., 2022, Watson et al., 2023), which we are currently exploring. For example, unconditional models can leverage inference-time guidance strategies (Wu et al., 2024), be fine-tuned conditionally (Denker et al., 2024) or in an amortized fashion for motif-scaffolding (Didi et al., 2023). For sequence conditioning and structure prediction, we can incorporate embeddings from language models (Penic et al., 2024, He et al., 2024).
@@ 3 Experiments
Evaluation metrics. We evaluate our models for unconditional RNA backbone generation, analogous to recent work in protein design (Yim et al., 2023b, a, Bose et al., 2023, Lin and AlQuraishi, 2023). We generate 50 backbones for target lengths sampled between 40 and 150 at intervals of 10. We then compute the following indicators of quality for these backbones:
  • Validity (scTM $\geq 0.45$): We inverse fold each generated backbone using gRNAde (Joshi et al., 2023) and pass $N_{\text{seq}}=8$ generated sequences into RhoFold (Shen et al., 2022). We then compute the self-consistency TM-score (scTM) between the predicted RhoFold structure and our backbone at the $C4^{\prime}$ level. We say a backbone is valid if scTM $\geq 0.45$; this threshold corresponds to roughly the same fold between two RNA strands (Zhang et al., 2022). We expand on this framework in Figure 2.
@@
Hyperparameters. We use 6 IPA blocks in our flow model, with an additional 3-layer torsion predictor MLP that takes in node embeddings from the IPA module. Our final model contains 16.8M trainable parameters. We use Adam optimizer with learning rate $0.0001$, $\beta_{1}=0.9$, $\beta_{2}=0.999$. We train for $120K$ gradient update steps on four NVIDIA GeForce RTX 3090 GPUs for about 15 hours with a batch size $B=20$. Each batch contains samples of the same sequence length to avoid padding. Further hyperparameters are listed in Appendix B.1.
4 Results

### v3

@@ Abstract
We introduce RNA-FrameFlow, the first generative model for 3D RNA backbone design. We build upon $SE(3)$ flow matching for protein backbone generation and establish protocols for data preparation and evaluation to address unique challenges posed by RNA modeling. We formulate RNA structures as a set of rigid-body frames and associated loss functions which account for larger, more conformationally flexible RNA backbones (13 atoms per nucleotide) vs. proteins (4 atoms per residue). Toward tackling the lack of diversity in 3D RNA datasets, we explore training with structural clustering and cropping augmentations. Additionally, we define a suite of evaluation metrics to measure whether the generated RNA structures are globally self-consistent (via inverse folding followed by forward folding) and locally recover RNA-specific structural descriptors. The most performant version of RNA-FrameFlow generates locally realistic RNA backbones of 40-150 nucleotides, over 40% of which pass our validity criteria as measured by a self-consistency TM-score $\geq 0.45$, at which two RNAs have the same global fold. Open-source code: github.com/rish-16/rna-backbone-design
@@ 1 Introduction
Designing RNA structures. Proteins, and the diverse structures they can adopt, drive essential biological functions in cells. Deep learning has led to breakthroughs in structural modeling and design of proteins (Jumper et al., 2021; Dauparas et al., 2022; Watson et al., 2023), driven by the abundance of 3D data from the Protein Data Bank (PDB). Concurrently, there has been a surge of interest in Ribonucleic Acids (RNA) and RNA-based therapeutics for gene editing, gene silencing, and vaccines (Doudna & Charpentier, 2014; Metkar et al., 2024). RNAs play several roles in cells: they carry genetic information coding for proteins (mRNA) as well as perform functions driven by their tertiary structural interactions (riboswitches and ribozymes). While there is growing interest in designing such structured RNAs for a range of applications in biotechnology and medicine (Mulhbacher et al., 2010; Damase et al., 2021), the current toolkit for 3D RNA design uses classical algorithms and heuristics to assemble RNA motifs as building blocks (Han et al., 2017; Yesselman et al., 2019). However, hand-crafted heuristics are not always broadly applicable across multiple tasks and rigid motifs may not fully capture the conformational dynamics that govern RNA functionality (Ganser et al., 2019; Li et al., 2023b). This presents an opportunity for deep generative models to learn data-driven design principles from existing 3D RNA structures.
@@
Figure 1: The RNA-FrameFlow pipeline for 3D backbone generation. Our implementation establishes RNA-specific protocols for data preparation and evaluation for FrameFlow (Yim et al., 2023a). (1) Each nucleotide in the RNA backbone is converted into a frame to parameterize the placement of $C4^{\prime}$ by a translation, $C3^{\prime}-C4^{\prime}-O4^{\prime}$ by a rotation, and the rest of the atoms via 8 torsion angles $\Phi$. (2) We train generative models on all RNA structures of length 40-150 nucleotides from RNAsolo (Adamczyk et al., 2022). We also explore training with structural clustering and cropping augmentations to tackle the lack of diversity in 3D RNA datasets. (3) We introduce evaluation metrics to measure the recovery of local structural descriptors and global self-consistency of designed structures via inverse-folding with gRNAde (Joshi et al., 2025) followed by forward-folding with RhoFold (Shen et al., 2022).
@@
What makes deep learning for RNA hard? The primary challenge is the paucity of raw 3D RNA structural data, manifesting as an absence of ML-ready datasets for model development (Joshi et al., 2025). Protein structure is primarily driven by hydrogen bonding along the backbone, and current deep learning models incorporate this inductive bias through backbone frames to represent residues (Jumper et al., 2021). RNA structure, however, is often more conformationally flexible and driven by base pairing interactions across strands as well as base stacking between rings of adjacent nucleotides (Vicens & Kieft, 2022).
@@
Additionally, RNA nucleotides, the equivalent of amino acids in proteins, include significantly more atoms as part of the backbone (13 compared to 4) which necessitates a generalization of backbone frames where the placement of most atoms needs to be parameterized by torsion angles. These complexities have contributed to relatively poor performance of deep learning for RNA structure prediction compared to proteins (Kretsch et al., 2023; Abramson et al., 2024). Additionally, structure prediction models cannot directly be used for designing or generating novel RNA structures with desired constraints. Thus, the lack of de novo RNA design models presents a gap worth addressing, which we aim to do in this work.
Our Contributions. We develop RNA-FrameFlow, the first generative model for 3D all-atom RNA backbone design, which trains on RNA of length 40-150 from the PDB; see Figure 1. We devise RNA frames to capture the significantly larger RNA nucleotides while still supporting all-atom modelling. Our method can unconditionally sample plausible backbones with over 40% validity, as measured by our custom evaluation pipeline that computes global self-consistency and local structural measurements. We also explore data augmentation protocols to address the paucity of diverse RNA structural data, which improves the novelty of designed RNA. We hope this work stimulates future research in generative models for RNA design.
@@ 2 The RNA-FrameFlow Pipeline
We are concerned with building a generative model that unconditionally outputs all-atom 3D RNA backbone structures, sampled from a distribution of realistic structures. Formally, given an RNA sequence of $N_{\text{nt}}$ nucleotides, we aim to generate a real-valued tensor $\mathbf{X}$ of shape ${N_{\text{nt}}\times 13\times 3}$ representing 3D coordinates for 13 backbone atoms per nucleotide. In the following sections, we will describe how we adapt FrameFlow (Yim et al., 2023a), an $SE(3)$ equivariant flow matching model for protein backbones, to our setting.
@@ 2.1 Representing RNA Backbones as Frames
Rationale for RNA frames. In proteins, it is standard to represent each residue by a frame centered at $C_{\alpha}$ with vectors along $C_{\alpha}-N$ and $C_{\alpha}-C$, and $O$ is placed assuming an idealized planar geometry (Jumper et al., 2021). However, no such canonical frame representation exists for RNAs. Furthermore, the RNA nucleotide contains significantly more backbone atoms (13) compared to proteins (4). As shown in Figure 1, the backbone fragment of RNA nucleotides comprise a phosphate group ($P,OP1,OP2,O5^{\prime}$), a ribose sugar ($C1^{\prime}-C5^{\prime},O2^{\prime},O3^{\prime},O4^{\prime}$), and a nitrogen atom $N$ at the stem of the base. To avoid the high modelling complexity of explicitly generating coordinates for all 13 backbone atoms, we decide to represent the atoms within each nucleotide as a rigid-body frame; this enables inferring the positions of all intra-nucleotide atoms via a frame center and orientation.
Constructing RNA frames. We select the $C4^{\prime},C3^{\prime}$, and $O4^{\prime}$ atoms to create the frame for each nucleotide, as in Morehead et al. (2023). All other backbone atoms are inferred with 8 torsions $\Phi=\{\phi_{i}\}_{i=1}^{8},\ \phi_{i}\in SO(2)$ that are predicted post-hoc after frame generation. The Gram-Schmidt process is used on $v_{1},v_{2}$ defined by the vectors along the $C4^{\prime}-O4^{\prime}$ and $C4^{\prime}-C3^{\prime}$ bonds; $C5^{\prime}$ is imputed based the positions of the other 3 atoms and tetrahedral geometry. Given the 8 torsion angles, we autoregressively place non-frame atoms in order of the torsions $\Phi$ in Figure 1, constructing the final set of all-atom RNA nucleotides. We describe this imputation of non-frame atoms as well as the choice of torsion angle parameterization in Appendix A.4.
Choice of frame atoms. We had two main considerations for selecting the atoms to create RNA frames: (1) the atoms should have roughly the same spatial orientation w.r.t. each other; and (2) the atoms should be reasonably close to the centroid in the nucleotide to reduce error accumulation when placing the furthest non-frame atoms. We choose $\{C3^{\prime},C4^{\prime},O4^{\prime}\}$ as these atoms spatially shift the least in naturally occurring RNA (Harvey & Prabhakaran, 1986). The non-frame backbone atoms – the remaining atoms in the ribose ring and the phosphate group atoms – are parameterized by torsion angles to account for their relative conformational flexibility. This choice of frame enables models to learn ring puckering, the planar rotation of the ribose ring about the $C4^{\prime}-C5^{\prime}$ bond which affects how the RNA interacts with partners to form complexes (Clay et al., 2017). We are actively evaluating alternate choices of RNA frames.
@@ 2.2 $SE(3)$ Flow Matching on RNA Frames
Input. Given a set of 3D coordinates, a simultaneous rotation and translation $(r,x)$ forms an orientation-preserving rigid-body transformation of the coordinates. The set of all such transformations in 3D is the Special Euclidean group $SE(3)$, which composes the group of 3D rotations $SO(3)$ and 3D translations in $\mathbb{R}^{3}$. We can represent an RNA frame $T=(r,x)$ as a translation $x\in\mathbb{R}^{3}$ from the global origin to place $C4^{\prime}$ and a rotation $r\in SO(3)$ to orient $C3^{\prime}-C4^{\prime}-O4^{\prime}$. Compared to working with raw 3D coordinates for each backbone atom, using the frame representation entails performing flow matching on the space of $SE(3)$. This is an inductive bias to reduce the degrees of freedom the generative model needs to learn. Instead of predicting 13 correlated 3D coordinates independently (39 quantities) for each nucleotide, we instead predict one 3D coordinate (of $C4^{\prime}$) and one 3$\times$3 rotation matrix (12 quantities). We follow Chen & Lipman (2024) and Yim et al. (2023a)’s framework for flow matching on $SE(3)$, which we summarise subsequently.
@@
Suppose $p_{0}(T_{0})$ and $p_{1}(T_{1})$ are the marginal distributions of randomly oriented and ground truth frames from our dataset of RNA structures, respectively. Suppose a time-dependent vector field $u_{t}$ leads to an ODE between the two distributions $p_{0}$ and $p_{1}$, i.e., assume there is a way to map from noisy samples to the corresponding true samples. This solution forms a ground truth probability path $p_{t}$ between the two distributions at time $t\in[0,1]$, which we can use to transform noisy samples to the true distribution. The continuity equation $\frac{\partial p}{\partial t}=-\nabla\cdot(p_{t}u_{t})$ relates the vector field $u_{t}$ to the evolution of the probability path $p_{t}$.
Given a noisy frame $T_{0}$ sampled from $p_{0}(T_{0})$ and the corresponding ground truth frame $T_{1}$ sampled from $p_{1}(T_{1})$, we construct a flow $T_{t}$ by following the probability path $p_{t}$ between $T_{0}$ and $T_{1}$ for any time step $t$ sampled from $\mathcal{U}(0,1)$. As shown by Chen & Lipman (2024) for the $SE(3)$ group (and other manifolds), the geodesic between the states $T_{0}$ and $T_{1}$ can be used to define an interpolation:
@@
The architecture of the flow model $v_{\theta}$ is similar to the structure module from AlphaFold2 comprising Invariant Point Attention layers interleaved with standard Transformer encoder layers, following Yim et al. (2023a; b). We use an MLP head to predict torsion angles $\Phi$.
@@
Suppose $S=\{C4^{\prime},C3^{\prime},O4^{\prime}\}$ is the set of frame atoms [^1] In Appendix B.1, we show how including more backbone atoms better accounts for larger RNA nucleotides and improves validity of generated samples. and the sequence length is $N$. We summarise the auxiliary losses subsequently.
@@
Conditional generation. The unconditional sampling strategy described above aims to generate realistic RNA backbones sampled from the training distribution. However using generative models in real-world design tasks entails conditional generation based on specified design constraints or requirements (Ingraham et al., 2022; Watson et al., 2023), which we are currently exploring. For example, unconditional models can leverage inference-time guidance strategies (Wu et al., 2024), be fine-tuned conditionally (Denker et al., 2024) or in an amortized fashion for motif-scaffolding (Didi et al., 2023). For sequence conditioning and structure prediction, embeddings from language models can also be incorporated (Penic et al., 2024; He et al., 2024).
@@ 3 Experiments
Evaluation metrics. We evaluate our models for unconditional RNA backbone generation, analogous to recent work in protein design (Yim et al., 2023b; a; Bose et al., 2023; Lin & AlQuraishi, 2023); see Figure 2. We generate 50 backbones for target lengths sampled between 40 and 150 at intervals of 10. We then compute the following indicators of quality for these backbones:
  • Validity (scTM $\geq 0.45$): We inverse fold each generated backbone using gRNAde (Joshi et al., 2025) and pass $N_{\text{seq}}=8$ generated sequences into RhoFold (Shen et al., 2022). We then compute the self-consistency TM-score (scTM) between the predicted RhoFold structure and our backbone at the $C4^{\prime}$ level. We say a backbone is valid if scTM $\geq 0.45$; this threshold corresponds to roughly the same fold between two RNAs (Zhang et al., 2022). Alternatively, we use an RMSD threshold of 4.3 Å, corresponding to the median RMSD of RhoFold on RNAsolo sequences.
@@
Hyperparameters. We use 6 IPA blocks in our flow model, with an additional 3-layer torsion predictor MLP that takes in node embeddings from the IPA module. Our final model contains 16.8M trainable parameters. We use AdamW optimizer with learning rate $0.0001$, $\beta_{1}=0.9$, $\beta_{2}=0.999$. We train for $120K$ gradient update steps on four NVIDIA GeForce RTX 3090 GPUs for $\sim$18 hours with a batch size $B=28$. Each batch contains samples of the same sequence length to avoid padding. Further hyperparameters are listed in Appendix A.1.
4 Backbone Generation Results